OpenStudy (vheah):

Linear algebra question. How do you sketch this Null space? (t/3, t, 0) ?

Hayhayz (hayhayz):

hi, can you state the definition of the null space

OpenStudy (vheah):

The set \[\left\{ x \epsilon R^n|Ax = 0\right\}\] 0 is 0 the vector and x is an x vector. @Hayhayz

OpenStudy (vheah):

it is also called the kernel of A when given A as an mxn matrix. But i dont quite understand it @Hayhayz

Hayhayz (hayhayz):

so okay for example for Ax=0 the null space is all the vectors x succh that this Ax=0 is true

OpenStudy (vheah):

Okay so what would be sample vectors x that would make that statement true?

Hayhayz (hayhayz):

yup

Hayhayz (hayhayz):

okay so in this case it is what it is

Hayhayz (hayhayz):

(t/3, t, 0) , all the vectors of this form seem to solve your matrix Ax=0

Hayhayz (hayhayz):

x=t/3,y=t,z=0

Hayhayz (hayhayz):